# Online Makespan Scheduling under Scenarios

### v1

@@ 4 Tight Bounds for Sufficiently Many Scenarios
A straightforward application of the pigeonhole principle yields that, if we have seven subhypergraphs $\mathcal{S}(1),\ldots,\mathcal{S}(7)$ that are each partitioned onto two online colors, at least three of them (without loss of generality $\mathcal{S}(1)$, $\mathcal{S}(2)$, $\mathcal{S}(3)$) must be partitioned onto the same two online colors. Again without loss of generality, these two online colors are $1$ and $2$. Considering Proposition 4, we can permute the colorings of the three copies $\mathcal{S}(1)$, $\mathcal{S}(2)$ and $\mathcal{S}(3)$ each offline so that for any offline color $u\in\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$, there exists a copy $\mathcal{S}(t)$ with two (online) monochromatic hyperedges $e_{i}^{t}\in E(\mathcal{S}(t))$ $(i\in\{1,2\})$ of size $2$ avoiding the color $u$. The hypergraph in Figure 4 restricted to nodes of the form $j_{s}^{i}$ illustrate these edges.
Once the six edges in question are found, one can insert nodes $x_{1}$, $x_{2}$, $x_{3}$ in this order, revealing hyperedges as depicted in Figure 4. To avoid a monochromatic edge, both $x_{1}$ and $x_{2}$ must be assigned the third online color. Keeping this in mind, it follows that $x_{3}$ cannot be assigned an online color without creating a monochromatic edge with respect to the online coloring. Together with the fact that the offline coloring indeed has makespan $1$, we obtain:
@@
Figure 4: All fifteen relevant nodes together with the hyperedges within them. The heights of the vertices encode their online color ($J_{1}$, $J_{2}$, $J_{3}$), while their printed colors denote their offline color. Any assignment of the node $x_{3}$, given the other nodes and induced hyperedges, leads to a makespan of $3$.
Subhypergraph $\mathcal{S}$ on Three Online Colors: The Palettes.
@@
For every online coloring $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$ of the nodes $V(\mathcal{S})=\{v_{1},\ldots,v_{7}\}$ for nonempty sets $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$, there is an offline coloring $c\colon V(\mathcal{S})\to\{\mathrm{red},\mathrm{blue},\mathrm{yellow}\}$ such that two vertices exist with different online colors but the same offline color.
@@
In this case, the subhypergraph $\mathcal{S}(t)$ is called a $C$ palette (i.e., yellow palette, blue palette or red palette) and the corresponding nodes $v_{i}(t)$ are called palette nodes. In particular, $c(v_{i}(t))\neq C$ holds for the palette nodes of a $C$ palette. The choice of the name is motivated by the fact that, if we connect a new node $w$ to e.g. yellow palette nodes, we are still allowed to extend the offline coloring by $c(w)=\mathrm{yellow}$, while maintaining that the makespan of the offline coloring is $1$.
The Construction.
Below is a summary of our construction for the case where Lemma 4.2 is not applicable. Essentially, we create copies of the subhypergraph $\mathcal{S}$ which we connect with the existing hypergraph. The online color of the node $v$ is denoted by $\varphi(v)$ and its (current) offline color by $c(v)$.
@@
Notice that we have maintained an offline coloring of the nodes with makespan $1$ so far up to the very last node, and the last node $n$ neighbors only blue and yellow nodes by construction and therefore the offline assignment $c(n)=\mathrm{red}$ does not increase the makespan. In total, this yields that the offline optimal value is $1$. A careful enumeration of introduced hyperedges reveals that $233$ hyperedges suffice to create a monochromatic hyperedge for any of the cases demonstrated, which proves Theorem 4.1.
A Note on Theorem 1.2.
Here, we abstain from presenting a proof of Theorem 1.2. Our construction in this general case is also based on incrementally creating monochromatic edges of each online color. However, we are faced with two fundamental challenges:
@@
    The monochromatic hyperedges increase in size rather slowly, yet unless they have size of $m-1$, they are not strong enough to act as a palette for other hyperedges in the same sense as here.

### v2

@@ 4.2 Pigeonholing the two-color case
A straightforward application of the pigeonhole principle yields that, if we have seven subgraphs $\mathcal{S}(1),\ldots,\mathcal{S}(7)$ that are each partitioned onto two online colors, at least three of them (without loss of generality $\mathcal{S}(1)$, $\mathcal{S}(2)$, $\mathcal{S}(3)$) must be partitioned onto the same two online colors. Again without loss of generality, these two online colors are $1$ and $2$. Considering Proposition 5, we can color the three copies $\mathcal{S}(1)$, $\mathcal{S}(2)$ and $\mathcal{S}(3)$ so as to satisfy the following properties:
  • There exist two hyperedges $e_{i}^{1}=\{j_{i}^{1},j_{i}^{2}\}\in E(\mathcal{S}(1))$ for $i\in\{1,2\}$ with the property that $c(e_{i}^{1})=\{\mathrm{red,blue}\}$,
  • there exist two hyperedges $e_{i}^{2}=\{j_{i}^{3},j_{i}^{4}\}\in E(\mathcal{S}(2))$ for $i\in\{1,2\}$ with the property that $c(e_{i}^{1})=\{\mathrm{blue,yellow}\}$,
  • there exist two hyperedges $e_{i}^{3}=\{j_{i}^{5},j_{i}^{6}\}\in E(\mathcal{S}(3))$ for $i\in\{1,2\}$ with the property that $c(e_{i}^{3})=\{\mathrm{yellow,red}\}$.
Moving forward, we restrict our subgraph to one with the nodes $j_{i}^{t}$ for $i\in\{1,2\},t\in\{1,\ldots,6\}$ and the hyperedges $k_{i}^{t}$ for $i\in\{1,2\},t\in\{1,2,3\}$. The remaining nodes in $V(\mathcal{S}(1))\dot{\cup}V(\mathcal{S}(2))\dot{\cup}V(\mathcal{S}(3))$ will not be connected to other nodes anymore, nor will their offline coloring be altered. Thus, they can be ignored from now on.
$j_{2}^{1}$$j_{1}^{1}$$j_{2}^{2}$$j_{1}^{2}$$e_{1}^{1}$$e_{2}^{1}$$j_{2}^{3}$$j_{1}^{3}$$j_{2}^{4}$$j_{1}^{4}$$e_{1}^{2}$$e_{2}^{2}$$j_{2}^{5}$$j_{1}^{5}$$j_{2}^{6}$$j_{1}^{6}$$e_{1}^{3}$$e_{2}^{3}$$J_{1}$$J_{2}$
Figure 11: The nodes $j_{i}^{t}$.
Now we introduce the thirteenth node $x_{1}$, revealing the full extent of the hyperedges $e_{1}^{1}=\{j_{1}^{1},j_{1}^{2},x_{1}\}$ and $e_{2}^{1}=\{j_{2}^{1},j_{2}^{2},x_{1}\}$. Unless a competitive ratio of $3$ is attained, the node $x_{1}$ must be assigned to the third online color. Thereafter, the fourteenth node $x_{2}$ is revealed together with hyperedges $e_{1}^{2}=\{j_{1}^{3},j_{1}^{4},x_{2}\}$, $e_{2}^{2}=\{j_{2}^{3},j_{2}^{4},x_{2}\}$, $f_{1}\supseteq\{x_{1},x_{2}\}$. Analogously, the fourteenth node is assigned to the third online color as well. Notice that we can feasibly extend the offline coloring by $c(x_{1})=\mathrm{yellow}$ and $c(x_{2})=\mathrm{red}$, as depicted in Figure 12.
Finally, we would like to introduce the last node $x_{3}$ with the hyperedges $e_{1}^{3}=\{j_{1}^{5},j_{1}^{6},x_{3}\}$, $e_{2}^{3}=\{j_{2}^{5},j_{2}^{6},x_{3}\}$ and $f=\{x_{1},x_{2},x_{3}\}$. For each online color, there exists a hyperedge $e\in\{e_{1}^{3},e_{2}^{3},f\}$ with $x_{3}\in e$ that already attains a makespan of $2$. Therefore, the node $x_{3}$ cannot be placed without increasing the makespan to $3$. On the other hand, extending the coloring $c$ by $c(v_{n})=\mathrm{blue}$, as suggested by Figure 12, yields a feasible coloring of the nodes, suggesting that the offline partitioning given by the colors attains a makespan of $1$ in every scenario.
As a consequence, we obtain:
@@
Figure 12: All fifteen relevant nodes together with the hyperedges within them. Notice that any assignment of the node $v_{n}$ leads to a makespan of $3$.
4.3 Subgraph $\mathcal{S}$ on Three Online Colors: The Palettes
@@
For every online coloring $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$ of the nodes $V(\mathcal{S})=\{v_{1},\ldots,v_{7}\}$ for nonempty sets $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$, there is an offline coloring $c\colon V(\mathcal{S})\to\{\mathrm{red},\mathrm{blue},\mathrm{yellow}\}$ such that two nodes exist with different offline colors but the same online color.
Proof 4.6.
We fix the offline coloring in Figure 10, i.e., nodes $3,4$ are red, nodes $2,7$ are blue and nodes $1,5,6$ are yellow. The only case where the claim does not hold for the coloring is when each $J_{i}$ consists of nodes of precisely one offline color. In other words, up to a permutation of the colors, we have $J_{1}=c^{-1}(\{\mathrm{red}\})$, $J_{2}=c^{-1}(\{\mathrm{blue}\})$, $J_{3}=c^{-1}(\{\mathrm{yellow}\})$. In this case, we alter the offline coloring as follows: $1$ and $5$ are blue instead of yellow, and $2$ is yellow instead of blue. It is straightforward to check that this is a feasible coloring, see Figure 13.
33112255446677
Figure 13: The second case in the proof of Lemma 4.5. Names of hyperedges are left out for brevity.
@@
In this case, the subhypergraph $\mathcal{S}(t)$ is called a $C$ palette (i.e., yellow palette, blue palette or red palette) and the corresponding nodes $v_{i}(t)$ are called palette nodes. In particular, $c(v_{i}(t))\neq C$. The choice of the name is motivated by the fact that, if we connect a new node $w$ to e.g. yellow palette nodes, we are still allowed to extend the offline coloring by $c(w)=\mathrm{yellow}$, while maintaining that the makespan of the offline coloring is $1$.
4.4 The Construction
Below is a summary of our construction for the case where Lemma 4.4 is not applicable. Essentially, we create copies of the subgraph $\mathcal{S}$ which we consecutively connect with the existing hypergraph in varying ways. Below is a summary of our construction for the case where Lemma 4.4 is not applicable. Essentially, we create copies of the subhypergraph $\mathcal{S}$ which we connect with the existing hypergraph. The online color of the node $v$ is denoted by $\varphi(v)$ and its (current) offline color by $c(v)$.
@@ 4.4 The Construction
Figure 14: The introduction of the last node $v_{n}$ leads to a makespan of $3$. Only the seven relevant nodes are shown, although the entire hypergraph has both an online coloring and a feasible offline coloring.
Step (i) is straightforward. To achieve the desired colors of $\mathcal{S}(1),\ldots,\mathcal{S}(7)$ as in step (ii), we find out the color $C$ of each $\mathcal{S}(t)$ (as a palette), and postcompose the coloring $c|_{V(\mathcal{S}(t))}\colon V(\mathcal{S}(t))\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ with the transposition
$\tau\colon\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$
that swaps the color $C$ with the designated color of $\mathcal{S}(t)$ as stated in Step (ii).
Step (iii): the subinstance $\mathcal{A}$.
Here, it suffices to connect the yellow palette $\mathcal{S}(1)$ to a new node $w_{\mathcal{A}}$ with the hyperedges $e=\{w_{1},v_{i}(1)\}$ for $t\in\{1,2,3\}$. By the definition of the palette nodes $v_{i}(1)$, they are not yellow. Since the new node $w_{\mathcal{A}}$ is only connected to non-yellow nodes, we can extend the offline coloring by $c(w_{\mathcal{A}})=\mathrm{yellow}$, maintaining its feasibility. For the $i_{\mathcal{A}}$ which holds $w_{\mathcal{A}}\in J_{i_{\mathcal{A}}}$, we have $e=\{w_{\mathcal{A}},v_{i_{\mathcal{A}}}(1)\}\in E(G[J_{i}])$ in particular.
Step (iv): the subinstance $\mathcal{B}$.
In this stage of the construction, we use the three palettes $\mathcal{S}(2)$, $\mathcal{S}(3)$, $\mathcal{S}(4)$ with palette nodes $v_{i}(t)$ for $i\in\{1,2,3\}$ for each copy $t\in\{2,3,4\}$. Note that we are utilizing precisely one palette of each color.
Now we introduce three new nodes $w_{\mathcal{B}1}$, $w_{\mathcal{B}2}$ and $w_{\mathcal{B}3}$. These nodes are to be understood as candidates to the node $w_{\mathrm{B}}$ that we would like to define. They are connected to each other as well as to the three $v_{i}(t)$ of their respective palette, i.e., there exists the hyperedge $\{w_{\mathcal{B}1},w_{\mathcal{B}2},w_{\mathcal{B}3}\}\in E$ as well as hyperedges $\{w_{\mathcal{B}t},v_{i}(t+1)\}$ for $t\in\{1,2,3\}$ and $i\in\{1,2,3\}$. Accordingly, we may tentatively color the nodes as $c(w_{\mathcal{B}1})=\mathrm{yellow}$, $c(w_{\mathcal{B}2})=\mathrm{blue}$ and $c(w_{\mathcal{B}3})=\mathrm{red}$. See Figure 15 for a sketch of the introduced subhypergraph.
$v_{2}(2)$$v_{1}(2)$$v_{2}(3)$$v_{1}(3)$$v_{3}(3)$$v_{3}(2)$$v_{1}(4)$$v_{2}(4)$$v_{3}(4)$$\mathcal{S}(2)$$\mathcal{S}(3)$$\mathcal{S}(4)$$w_{\mathcal{B}1}$$w_{\mathcal{B}2}$$w_{\mathcal{B}3}$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$
Now we would like to choose the node $w_{\mathcal{B}}$ ultimately. To this end, we consider any online assignment of the three candidates $w_{\mathcal{B}t}$ ($t=1,2,3$). If $\varphi(w_{\mathcal{B}1})=\varphi(w_{\mathcal{B}2})=\varphi(w_{\mathcal{B}3})$, we obtain a makespan of $3$ and can immediately truncate the instance, as it no longer has a competitive ratio better than $3$. Hence, there are at least two online colors $i\in\{1,2,3\}$ that the candidate nodes $w_{\mathcal{B}t}$ are assigned to, and at least one of them is not $i_{A}$. We pick such an online color and call it $i_{B}$.
Step (v): the subinstance $\mathcal{C}$.
To ensure that an edge $e_{i^{\prime\prime}}\in E(G[J_{i^{\prime\prime}}])$ is created in precisely the online color $i^{\prime\prime}\notin\{i,i^{\prime}\}$, we introduce a number of candidate nodes $w_{\mathcal{C}s}$ similar to the previous step. Here, it is integral that the nodes $w_{\mathcal{C}s}$ are connected to each other in a way that is
  1. (i)
    dense enough so that they are forced to utilize all three online colors, but also
  2. (ii)
    sparse enough that the offline coloring can be extended to the $w_{\mathcal{C}s}$.
Conveniently, we can recollect to our considerations about the subgraph $\mathcal{S}$: If we create seven copies of it, one of them is forced to attain all three online colors, or else we can complete the hypergraph in a way that no online coloring is better than $3$-competitive.
Recalling that we may connect a node $w_{j}$ with the three nodes $v_{i}(t)$ of the palette in its offline color, we define the seven nodes as depicted in Figure 10. More precisely, we create palettes $\mathcal{S}(5)$, $\mathcal{S}(6)$, $\mathcal{S}(7)$ and introduce nodes $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$ together with the following hyperedges:
  • $\{w_{\mathcal{C}1},w_{\mathcal{C}2},w_{\mathcal{C}3}\}$, $\{w_{\mathcal{C}2},w_{\mathcal{C}4},w_{\mathcal{C}5}\}$, $\{w_{\mathcal{C}4},w_{\mathcal{C}6},w_{\mathcal{C}7}\}$, $\{w_{\mathcal{C}3},w_{\mathcal{C}6}\}$, $\{w_{\mathcal{C}3},w_{\mathcal{C}7}\}$, as well as their subsets of size $2$ (so that $G[\{w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}\}]\cong\mathcal{S}$),
  • $\{v_{i}(5),w_{\mathcal{C}s}\}$ for $i\in\{1,2,3\}$, $s\in\{1,5,6\}$, i.e., between yellow palette nodes and new nodes colored yellow in Figure 10,
  • $\{v_{i}(6),w_{\mathcal{C}s}\}$ for $i\in\{1,2,3\}$, $s\in\{2,7\}$, i.e., between blue palette nodes and new nodes colored blue in Figure 10,
  • $\{v_{i}(7),w_{\mathcal{C}s}\}$ for $i\in\{1,2,3\}$, $s\in\{3,4\}$, i.e., between red palette nodes and new nodes colored red in Figure 10.
An overview of the construction, without specifying the online colors of the nodes $w_{\mathcal{C}s}$, is given in Figure 16.
$v_{1}(1)$$v_{2}(1)$$v_{3}(1)$$v_{1}(2)$$v_{2}(2)$$v_{3}(2)$$v_{1}(3)$$v_{2}(3)$$v_{3}(3)$$w_{\mathcal{C}3}$$w_{\mathcal{C}1}$$w_{\mathcal{C}2}$$w_{\mathcal{C}5}$$w_{\mathcal{C}4}$$w_{\mathcal{C}6}$$w_{\mathcal{C}7}$$\mathcal{S}(5)$$\mathcal{S}(7)$$\mathcal{S}(6)$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$
@@
Recall that our goal is to find an edge of the form $\{v_{i^{\prime\prime}}(t),w_{\mathcal{C}s}\}\subseteq J_{i^{\prime\prime}}$ for the $i^{\prime\prime}\notin\{i,i^{\prime}\}$. If the nodes $w_{\mathcal{C}s}$ for $s\in\{1,\ldots,7\}$ are spread across the three online colors, there is an index $s\in\{1,\ldots,7\}$ with the property that $w_{\mathcal{C}s}\in J_{i^{\prime\prime}}$; for a suitable $t$ depending on the offline color of the node $w_{\mathcal{C}s}$, we thus find the node $v_{i^{\prime\prime}}(t)$ and hence the aforementioned edge as desired.
In order to guarantee the existence of $w_{\mathcal{C}s}\in J_{i^{\prime\prime}}$, we need to convince ourselves that the subgraph induced by $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$ is partitioned on all three online colors. Conveniently, we have seen in Subsection 4.2 that if we create seven copies of the subgraph $G[\{w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}\}]$, one of the copies must admit all three online colors or else the assignment cannot be better than $3$-competitive by Lemma 4.4. Therefore, we create seven copies of the subgraph induced by $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$, combining the nodes in each copy with the corresponding palette nodes.
In total, one of the seven copies of $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$ must admit all three online colors, in particular the online color $i^{\prime\prime}$. For a node $w_{\mathcal{C}s}$ attaining this online color, we find a palette node $v_{i^{\prime\prime}}(t)$ such that $\{v_{i^{\prime\prime}}(t),w_{\mathcal{C}s}\}\in J_{i^{\prime\prime}}$. Now we compose the offline coloring $c$ obtained so far with a permutation of the three offline colors such that $c(v_{i^{\prime\prime}}(t))=\mathrm{blue}$ and $c(w_{\mathcal{C}s})=\mathrm{yellow}$.
The reader might wonder why we need all three components; indeed, we will have three hyperedges of the type $e_{i}$ already in the underlying hypergraph of $\mathcal{C}$ alone. However, we require that the hyperedges $e_{i}$ are each colored in blue and yellow endpoints, which is not possible when we take all of them from $\mathcal{C}$.
Step (vi): Recoloring the connected components.
Finally, we permute the colors in the three connected components created in Steps (iii), (iv) and (v) (as shown in Figure 15) such that $c(w_{\mathcal{B}{i_{\mathcal{B}}}})=\mathrm{yellow}$ and $c(v_{i_{\mathcal{B}}}(i_{\mathcal{B}}))=\mathrm{blue}$. See Figure 17 for an example of the permutation process.
$v_{1}(4)$$v_{2}(4)$$v_{3}(4)$$\mathcal{S}(4)$$w_{\mathcal{B}3}$112233
Figure 17: Permuting the colors of the palette $\mathcal{S}(4)$. This permutation was implied by the assignment $\varphi(w_{\mathcal{B}3})=i_{\mathcal{B}}=3$. We have to ensure that $c(v_{3}(3))=\mathrm{blue}$ and $c(w_{\mathcal{B}3})=\mathrm{yellow}$, which is achieved by transposing the offline colors yellow and blue. Of course, this transposition is applied to the entire component, such that e.g. $c(w_{\mathcal{B}1})=\mathrm{blue}$, which is omitted here.
We are now ready to execute Step (vii) and force instance to admit a makespan of $3$. To this end, we summarize our setting so far.
Lemma 4.7.
There exists a family $\mathcal{I}$ of instances over the node set $V$ together with feasible colorings $c\colon V\to\{\mathrm{blue},\mathrm{red},\mathrm{yellow}\}$ such that for every online assignment $\varphi\colon V\to\{1,2,3\}$, at least one instance $I\in\mathcal{I}$ satisfies one of the following properties:
@@
    $\mathrm{makespan}(I)=3$, or